(A) Find the largest x-interval where the initial value problem has a unique solution:

(x^{2}-4x)y^{(6)}+\frac{1}{ln(x-1)}y^{(5)}+(x+1)y'''+e^{x}y' +(tanx)y = \frac{1}{x^{2}-9}; y(2.5)=A, y'(2.5)=B,y''(2.5)=C,y'''(2.5)=D,y^{(4)}(2.5)=E,y^{(5)}(2.5)=F,

Where A, B, C, D, E, F are some known constants.

(B) Determine whether the set of functions {5,sin^{2}x,cos2x} could form a fundamental set of solution of a linear differential equation

Thank you
Answer #1

We first state a hypothesis Hypothesis to consider the nth-order lineer differential equation Aota) dan tin Amlosg = F2-0 7 w*) let fica = 5 ; fzco) - Sinha; fz(n) = cos 22 . fila)= fi(a) = 0; fal () = 2.Sina cos 2 = Sin27 ; fè (m) =20052* , The wr